# Abundant Population of Broad H Emitters in the GOODS-N Field Revealed by CONGRESS, FRESCO, and JADES

### v1

@@ 4 Results
Among all broad H emitters, the broad H components exhibit average FWHM of 2050 km/s (ranging from 1077ā3550 km/s) and luminosity of 4.52 erg/s (ranging from 1.1ā32.9 erg/s). [NĀ II] emission with a S/N3 is detected in only two objects (i.e., GN-1034620 and GN-1089568). H absorption is also found to be rare within our sample, with a robust detection observed in only one object (i.e., GN-1087388), consistent with Matthee etĀ al. (2024), who also reported the presence of an absorption feature in this source. However, we cannot fully rule out the presence of H absorption features in some of these emitters. As noted earlier, recent studies with NIRSpec have shown that detecting such absorption features typically requires high S/N (e.g., Juodžbalis etĀ al. 2024; Ji etĀ al. 2025; DāEugenio etĀ al. 2025b; Rusakov etĀ al. 2025). Given the relatively low S/N of many sources in our sample, it remains possible that H absorption is present but remains undetected in some cases.
@@
Although the line centroids of the broad and narrow components are allowed to be different, the majority of objects exhibits offsets of 100 km/s, with the largest offset of ākm/s observed in GN-1086855. We find that the broad components contribute, on average, about 74% of the total H luminosity, with only one object exhibiting a broad component contribution of less than 50%. We note that our sample exhibits a higher average contribution from the broad component compared to that of the Type 1 AGN identified with NIRSpec in JADES (i.e., 46% 25%; Juodžbalis etĀ al. 2025). This difference is likely due to the lower S/N of our sample, which may hinder the detection of faint broad components. However, by taking the uncertainties into consideration, our average broad-component contribution remains consistent with that of Juodžbalis etĀ al. (2025) at the 1- level.
@@
Figure 2 presents the false-color stamp images of the 19 discovered broad H emitters while Figure 3 illustrates the sky positions of these emitters, overlaid with the footprints of CONGRESS, FRESCO, and JADES. Notably, our sample includes 7 broad H emitters from FRESCO, as reported by Matthee etĀ al. (2024), and 3 broad H emitters from CONGRESS, as reported by Covelo-Paz etĀ al. (2024). We note that two broad H emitters in Covelo-Paz etĀ al. (2024) were excluded as they do not meet our selection criteria. A detailed explanation is provided in AppendixĀ A.
@@
We highlight 9 broad H emitters newly identified in this work in Table 1 and 2. The exact reasons for their absence in previous studies remain unclear. For the seven CONGRESS sources not repoted in Covelo-Paz etĀ al. (2024), the paper does not provide enough information to make assessment. Regarding the two FRESCO sources not reported in Matthee etĀ al. (2024), we suspect that GN-1090549 may have been excluded due to its relatively low S/N, as Matthee etĀ al. (2024) adopted a stricter cut (S/N 5) than we did. Although GN-1090549 shows a broad component with S/N 5 in our analysis, differences in 1D grism spectra extraction methods between the two studies may explain this discrepancy. For GN-1034620, we note that this source overlaps with a giant local galaxy, and it is possible that segmentation differences led to it being excluded from their sample.
4.2 Narrow [OĀ III] Lines and AGN Origin of the Broad H Lines
@@
Since the JWST/NIRCam grism spectra in our study only cover two wideband filters (i.e., F356W and F444W), the [OĀ III] doublet is within the wavelength coverage for only three emitters in our sample. Figure 4 shows the 1D line fitting results for these three sources. Since the FWHM observed in [OĀ III] is significantly smaller than that observed in H, this effectively rules out the possibility that the broad components originate from galactic outflows.
@@
For emitters without [OĀ III] coverage, although we cannot entirely exclude the possibility of galactic outflows, the majority exhibit broad component FWHM values significantly larger than 1000 km/s, along with broad-to-narrow H flux ratios exceeding 50%. In contrast, broad components associated with galactic outflows typically exhibit narrower FWHMs, ranging from a few hundred to 1000 km/s, depending on whether the outflows are driven by star formation or AGN activity, and show broad-to-narrow H flux ratios of 50% (see Matthee etĀ al. 2024 and references therein). We also note that in published NIRSpec studies, nearly all broad H emitters show narrow [OĀ III] emission, with their broad Balmer lines consistently attributed to the AGN broad line region (BLR) (e.g., Harikane etĀ al. 2023; Greene etĀ al. 2024). Therefore, it is reasonable to assume that the broad components in our sample are primarily caused by AGN activity.
JADES IDaaIDs are from JADES DR2 (DāEugenio etĀ al., 2025a). | RA | DEC | FWHM | F356W | ||
@@
GN-1087315bbNew broad H emitters identified in this work. The other emitters are already reported by Covelo-Paz etĀ al. (2024). | 189.333584 | 62.246178 | 3.91 | 12.51.0 | 1514183 | 25.73 |
GN-1082263bbNew broad H emitters identified in this work. The other emitters are already reported by Covelo-Paz etĀ al. (2024). | 189.212584 | 62.227436 | 3.98 | 13.61.6 | 1195187 | 24.82 |
GN-1089568bbNew broad H emitters identified in this work. The other emitters are already reported by Covelo-Paz etĀ al. (2024). | 189.151821 | 62.272229 | 4.05 | 23.21.2 | 1462143 | 24.90 |
GN-1029154bbNew broad H emitters identified in this work. The other emitters are already reported by Covelo-Paz etĀ al. (2024). | 189.159025 | 62.260221 | 4.17 | 20.71.5 | 1932204 | 24.96 |
GN-1086784 | 189.305706 | 62.236946 | 4.40 | 10.60.8 | 2414263 | 25.42 |
GN-1008411bbNew broad H emitters identified in this work. The other emitters are already reported by Covelo-Paz etĀ al. (2024). | 189.211089 | 62.250271 | 4.41 | 9.41.6 | 3217692 | 25.44 |
GN-1086855 | 189.286512 | 62.238138 | 4.41 | 24.45.4 | 3550860 | 24.61 |
@@
GN-1033320bbNew broad H emitters identified in this work. The other emitters are already reported by Covelo-Paz etĀ al. (2024). | 189.125779 | 62.287404 | 4.48 | 8.21.1 | 1951398 | 25.54 |
GN-1085355bbNew broad H emitters identified in this work. The other emitters are already reported by Covelo-Paz etĀ al. (2024). | 189.094365 | 62.198974 | 4.88 | 10.70.8 | 1801158 | 25.42 |
JADES IDaaIDs are from JADES DR2 (DāEugenio etĀ al., 2025a). | RA | DEC | FWHM | FWHM | F444W | |||
@@
GN-1090253 | 189.285544 | 62.280781 | 5.09 | 10.00.7 | 1455105 | Ā | Ā | 24.65 |
GN-1014406 | 189.072090 | 62.273431 | 5.15 | 13.20.8 | 2313181 | Ā | Ā | 25.44 |
GN-1034620bbNew broad H emitters identified in this work. The other emitters are already reported by Matthee etĀ al. (2024). | 189.159764 | 62.295924 | 5.19 | 21.11.1 | 107771 | Ā | Ā | 24.31 |
GN-1090549bbNew broad H emitters identified in this work. The other emitters are already reported by Matthee etĀ al. (2024). | 189.235941 | 62.285544 | 5.20 | 7.51.0 | 1721307 | Ā | Ā | 25.92 |
GN-9994014 | 189.300125 | 62.212044 | 5.23 | 19.51.1 | 2084156 | Ā | Ā | 24.92 |
GN-1013188 | 189.057100 | 62.268940 | 5.25 | 8.40.6 | 2187163 | Ā | Ā | 25.38 |
@@
5 Discussion
5.1 The ā Relation
To gain insights into the accretion efficiency and growth rates of our broad H emitters, we estimated the black hole mass () and AGN bolometric luminosity () of our broad H emitters following the methodology widely adopted in recent studies (e.g., Ćbler etĀ al. 2023; Kocevski etĀ al. 2023; Harikane etĀ al. 2023; Matthee etĀ al. 2024). More specifically, are estimated using the virial relation proposed by Reines etĀ al. (2013),
where represents the scale factor, which we assume to be 1.075, consistent with the value adopted in Reines & Volonteri (2015).
Regarding AGN bolometric luminosity, we follow the approach described in Harikane etĀ al. (2023), estimating it from the luminosity of the broad H component (Greene & Ho, 2005) and applying the corresponding bolometric correction from Richards etĀ al. (2006):
@@ 5.1 The ā Relation
We plot the versus H-inferred AGN bolometric luminosities in Figure 5. For comparison, we include UV-selected luminous quasars at from Yang etĀ al. (2021), broad-line AGN (BLAGN) identified with NIRSpec grating spectroscopy at from Maiolino etĀ al. (2024)111We exclude three sources from Maiolino etĀ al. (2024) that exhibit two broad components., BLAGN at from Harikane etĀ al. (2023), and BLAGN detected with NIRSpec prism spectroscopy at from Greene etĀ al. (2024). Notably, the for those BLAGN were derived using the same methodology employed in this study, whereas Yang etĀ al. (2021) estimated based on the continuum luminosity at 3000 Ć Ā and the FWHM of the MgĀ II line.
@@
The majority of our sample is found to be accreting at an Eddington ratio of 0.1, lower than that of UV-selected luminous quasars, consistent with the intrinsically faint nature of our broad H emitters. These moderate accretion rates are in line with those reported for BLAGN in Maiolino etĀ al. (2024) and Greene etĀ al. (2024). In contrast, the bolometric luminosities of the BLAGN in Harikane etĀ al. (2023) are derived using a more complex methodology, where as estimated from EquationĀ 5 is treated as a lower limit. Taking this into account, we find no significant difference in the inferred accretion rates between our sample and that of Harikane etĀ al. (2023).
@@
Interestingly, we notice that the BLAGN identified with NIRSpec prism() spectroscopy in Greene etĀ al. (2024) tend to occupy the high-mass end of the distribution relative to our sample. The recent JWST Cycle 3 General Observer program Slitless Areal Pure-Parallel High-Redshift Emission Survey (SAPPHIRES-PID: 6434; PI: E. Egami; Sun etĀ al. 2025a) observed one BLAGN (FWHMāā4000ākm/s) from Greene etĀ al. (2024) with the F444W grism, providing an independent H line measurement. While the source still exhibits a broad feature, the line width appears significantly narrower (FWHMā1300ākm/s). Given the grismās higher spectral resolution, we adopt the latter measurement and mark the resulting and with an āXā in FigureĀ 5, which suggests the possibility that some of the line widths measured by NIRSpec/prism may be overestimated. The details about the analysis of the SAPPHIRES spectrum are presented in AppendixĀ B
5.2 The - relation
@@
In this work, the stellar masses () of the AGN hosts are measured by SED-fitting using the Prospector Bayesian inference framework (Leja etĀ al. 2017; Johnson etĀ al. 2021, 2022) based on the Flexible Stellar Population Synthesis (FSPS; Conroy etĀ al. 2009; Conroy & Gunn 2010) code. To achieve reliable AGN identification, we adopt a modified version of Prospector presented in Lyu etĀ al. (2024), which incorporates a semi-empirical model for AGN UV-to-mid-IR continua with nebular emission lines and dust attenuation and replaces the default AGN torus model in Prospector.
Examples of galaxy SED-fitting results are presented in Figure 6. Due to the limited wavelength coverage, we have only seen AGN evidence from SED analysis in 10 out of 19 galaxies despite the existence of a broad H emission line. Therefore, follow-up MIRI observations to obtain mid-infrared data for these broad H emitters would be valuable for better constraining their physical properties. Table 3 lists the best-fit values of for all broad H emitters. The typical statistical uncertainty in estimates derived using Prospector is less than 0.1ā0.15 dex, which is smaller than the systematic uncertainties introduced by variations (e.g., 0.2ā0.3 dex) in SED model choices, such as the star formation history. Therefore, we adopt a conservative uncertainty of 0.3 dex for our measurements. Table 3 also summarizes the key physical properties of the 19 broad H emitters.
@@
Figure 7 illustrates the distribution of broad H emitters identified in this study on the plane. To further explore the evolution of the relation over cosmic time, we incorporate AGN recently discovered with JWST across a wide redshift range. These include one potential AGN at , UHZ1 (BogdÔn et al., 2024), as well as BLAGN in the GOODS-N field at (Maiolino et al., 2024) and (Sun et al., 2025c), and BLAGN at identified by Harikane et al. (2023). The local relations for AGN hosts derived in Reines & Volonteri (2015) (hereafter RV15) and Greene et al. (2020) (all, limits case, hereafter GSH20) are also included for comparison purposes.
@@
We find that nearly all AGN recently revealed by JWST observations at lie above RV15, with the majority located significantly above the relation. When compared to the GSH20, which is considered a more robust local relation that shows no bias with galaxy type and provides a better match to broad-line selected AGN at (Sun etĀ al., 2025b), the offset is somewhat reduced. Nevertheless, many sources still exhibit higher / ratios than those predicted by GSH20. The over-massive relative to the of host galaxy, when compared with the local scaling relations, is consistent with findings from numerous previous studies of AGN at similar redshifts (e.g., Kocevski etĀ al. 2023; Ćbler etĀ al. 2023). These results may suggest a different evolution path of the ā relation in the early Universe compared to that observed locally. Additionally, the presence of SMBHs with above the local scaling relations aligns with theoretical predictions from models invoking heavy black hole seeds and super-Eddington accretion, which are proposed to explain the rapid growth of SMBHs at very early time (Trinca etĀ al., 2022; Volonteri etĀ al., 2023; Schneider etĀ al., 2023). Alternatively, the growth of the host galaxy can also be suppressed by strong feedback processes (see Zhu etĀ al. 2025 and references therein). However, it is also important to note that, although our AGN sample is selected based on the presence of broad emission lines in spectroscopy, selection bias still exists to some extent. Specifically, due to the detection limits of the grism observations and the requirement of FWHM 1000 km/s, we may fail to detect very faint, low-mass AGN (see Section 5.4 for more details), which could otherwise help fill the observed gap between our sample and the local - relation.
5.3 Are Broad Line Emitters LRDs?

### v2

@@ IV Results
Among all broad H emitters, the broad H components exhibit average FWHM of 2032 km/s (ranging from 1077–3281 km/s) and luminosity of 4.40 erg/s (ranging from 0.9–32.9 erg/s). [N II] emission with a S/N3 is detected in only two objects (i.e., GN-1034620 and GN-1089568). H absorption is also found to be rare within our sample, with a robust detection observed in only one object (i.e., GN-1087388), consistent with Matthee et al. (2024), who also reported the presence of an absorption feature in this source. However, we cannot assess the presence of H absorption with grism data. As noted earlier, recent studies with NIRSpec have shown that detecting such absorption features typically requires high S/N or higher resolution (e.g., Juodžbalis et al. 2024; Ji et al. 2025; D’Eugenio et al. 2025b; Rusakov et al. 2025). Given the relatively low S/N of many sources in our sample, it remains possible that H absorption is present but remains undetected in some cases.
@@
Although the line centroids of the broad and narrow components are allowed to be different, the majority of objects exhibits offsets of 100 km/s, with the largest offset of km/s observed in GN-1086855. We find that the broad components contribute, on average, about 72% of the total H luminosity, with only one object exhibiting a broad component contribution of less than 50%. We note that our sample exhibits a higher average contribution from the broad component compared to that of the Type 1 AGN identified with NIRSpec in JADES (i.e., 46% 25%; Juodžbalis et al. 2025). This difference is likely due to the lower S/N of our sample, which may hinder the detection of faint broad components. However, by taking the uncertainties into consideration, our average broad-component contribution remains consistent with that of Juodžbalis et al. (2025) at the 1- level.
@@
Figure 2 presents the false-color stamp images of the 19 discovered broad H emitters while Figure 3 illustrates the sky positions of these emitters, overlaid with the footprints of CONGRESS, FRESCO, and JADES. Notably, our sample includes 7 broad H emitters from FRESCO, as reported by Matthee et al. (2024), and 3 broad H emitters from CONGRESS, as reported by Covelo-Paz et al. (2024). We note that two broad H emitters in Covelo-Paz et al. (2024) were excluded as they do not meet our selection criteria. A detailed explanation is provided in Appendix A.
@@
We highlight 9 broad H emitters newly identified in this work in Table 1 and 2. The exact reasons for their absence in previous studies remain unclear. For the seven CONGRESS sources not repoted in Covelo-Paz et al. (2024), the paper does not provide enough information to make assessment. Regarding the two FRESCO sources not reported in Matthee et al. (2024), we suspect that GN-1090549 may have been excluded due to its relatively low S/N, as Matthee et al. (2024) adopted a stricter cut (S/N 5) than we did. Although GN-1090549 shows a broad component with S/N 5 in our analysis, differences in 1D grism spectra extraction methods between the two studies may explain this discrepancy. For GN-1034620, we note that this source overlaps with a giant local galaxy, and it is possible that segmentation differences led to it being excluded from their sample.
IV.2 Narrow [O III] Lines and AGN Origin of the Broad H Lines
@@
Since the JWST/NIRCam grism spectra in our study only cover two wideband filters (i.e., F356W and F444W), the [O III] doublet is within the wavelength coverage for only three emitters in our sample. Figure 4 shows the 1D line fitting results for these three sources. Since the FWHM observed in [O III] is significantly smaller than that observed in H, this effectively rules out the possibility that the broad components originate from galactic outflows.
@@
For emitters without [O III] coverage, although we cannot entirely exclude the possibility of galactic outflows, the majority exhibit broad component FWHM values significantly larger than 1000 km/s, along with broad-to-narrow H flux ratios exceeding 50%. In contrast, broad components associated with galactic outflows typically exhibit narrower FWHMs, ranging from a few hundred to 1000 km/s, depending on whether the outflows are driven by star formation or AGN activity, and show broad-to-narrow H flux ratios of 50% (see Matthee et al. 2024 and references therein). We also note that in published NIRSpec studies, nearly all broad H emitters show narrow [O III] emission, with their broad Balmer lines consistently attributed to the AGN broad line region (BLR) (e.g., Harikane et al. 2023; Greene et al. 2024). Therefore, it is reasonable to assume that the broad components in our sample are primarily caused by AGN activity.
| JADES IDaaIDs are from JADES DR2 (D’Eugenio et al., 2025a). | RA | DEC | FWHM | F356W | ||
@@
| GN-1087315bbNew broad H emitters identified in this work. The other emitters are already reported by Covelo-Paz et al. (2024). | 189.333584 | 62.246178 | 3.91 | 12.51.0 | 1514183 | 25.73 |
| GN-1082263bbNew broad H emitters identified in this work. The other emitters are already reported by Covelo-Paz et al. (2024). | 189.212584 | 62.227436 | 3.98 | 9.61.3 | 1083205 | 24.82 |
| GN-1089568bbNew broad H emitters identified in this work. The other emitters are already reported by Covelo-Paz et al. (2024). | 189.151821 | 62.272229 | 4.05 | 23.21.2 | 1462143 | 24.90 |
| GN-1029154bbNew broad H emitters identified in this work. The other emitters are already reported by Covelo-Paz et al. (2024). | 189.159025 | 62.260221 | 4.17 | 20.81.5 | 2003225 | 24.96 |
| GN-1086784 | 189.305706 | 62.236946 | 4.40 | 11.71.2 | 3179504 | 25.42 |
| GN-1008411bbNew broad H emitters identified in this work. The other emitters are already reported by Covelo-Paz et al. (2024). | 189.211089 | 62.250271 | 4.41 | 9.41.7 | 3281757 | 25.44 |
| GN-1086855 | 189.286512 | 62.238138 | 4.41 | 16.92.3 | 1724271 | 24.61 |
@@
| GN-1033320bbNew broad H emitters identified in this work. The other emitters are already reported by Covelo-Paz et al. (2024). | 189.125779 | 62.287404 | 4.48 | 8.21.1 | 1951398 | 25.54 |
| GN-1085355bbNew broad H emitters identified in this work. The other emitters are already reported by Covelo-Paz et al. (2024). | 189.094365 | 62.198974 | 4.88 | 10.70.8 | 1801158 | 25.42 |
| JADES IDaaIDs are from JADES DR2 (D’Eugenio et al., 2025a). | RA | DEC | FWHM | FWHM | F444W | |||
@@
| GN-1090253 | 189.285544 | 62.280781 | 5.09 | 10.00.7 | 1455105 | 24.65 | ||
| GN-1014406 | 189.072090 | 62.273431 | 5.15 | 13.21.5 | 3212639 | 25.44 | ||
| GN-1034620bbNew broad H emitters identified in this work. The other emitters are already reported by Matthee et al. (2024). | 189.159764 | 62.295924 | 5.19 | 21.11.1 | 107771 | 24.31 | ||
| GN-1090549bbNew broad H emitters identified in this work. The other emitters are already reported by Matthee et al. (2024). | 189.235941 | 62.285544 | 5.20 | 7.51.0 | 1721307 | 25.92 | ||
| GN-9994014 | 189.300125 | 62.212044 | 5.23 | 19.51.1 | 2084156 | 24.92 | ||
| GN-1013188 | 189.057100 | 62.268940 | 5.25 | 8.00.6 | 1957147 | 25.38 | ||
@@
V Discussion
V.1 Luminosity function of broad-line H emitters
We compute the broad H luminosity function (LF) for the 19 broad H emitters identified in this work. The broad H LF is computed following the direct method (Schmidt, 1968):
where is the broad H luminosity of each LF bin, and represent the completeness correction and the maximum survey volume for the i-th broad H emitter, respectively. The completeness correction and maximum survey volume are measured following the methodology described in Lin et al. (2024). The uncertainties in the LF is estimated via Monte Carlo experiments. Table 3 summarizes the broad H LF measured in this work.
We present the derived broad H LF of our sample in Figure 5. For comparison, we also include the broad H LFs derived from JWST-selected broad-line AGN in other studies (Matthee et al., 2024; Lin et al., 2024, 2025d), as well as the quasar broad H LF at , obtained by converting the bolometric LFs from Shen et al. (2020). The conversion assumes the empirical relations between the bolometric luminosity and the AGN-induced H luminosity from Greene and Ho (2005), with the bolometric correction from Richards et al. (2006). We find that at the faint end (i.e., ), the number density of broad H emitters is significantly higher than the extrapolation of the quasar LF. At the bright end, however, the LF begins to converge toward that of quasars. This trend is consistent with the LFs derived from other JWST-selected AGN in the literature, suggesting a prevalence of broad-line AGN in the early universe.
| N | |||
| 42.117 | 0.414 | 7 | |
| 42.531 | 0.414 | 10 | |
| 43.153 | 0.819 | 2 |
V.2 The – Relation
To gain insights into the accretion efficiency and growth rates of our broad H emitters, we estimated the black hole mass () and AGN bolometric luminosity () of our broad H emitters following the methodology widely adopted in recent studies (e.g., Übler et al. 2023; Kocevski et al. 2023; Harikane et al. 2023; Matthee et al. 2024). More specifically, are estimated using the virial relation proposed by Reines et al. (2013),
where represents the scale factor, which we assume to be 1.075, consistent with the value adopted in Reines and Volonteri (2015).
Regarding AGN bolometric luminosity, we follow the approach described in Harikane et al. (2023), estimating it from the luminosity of the broad H component (Greene and Ho, 2005) and applying the corresponding bolometric correction from Richards et al. (2006):
@@ V.2 The – Relation
We plot the versus H-inferred AGN bolometric luminosities in Figure 6. For comparison, we include UV-selected luminous quasars at from Yang et al. (2021), broad-line AGN (BLAGN) identified with NIRSpec grating spectroscopy at from Maiolino et al. (2024)111We exclude three sources from Maiolino et al. (2024) that exhibit two broad components., BLAGN at from Harikane et al. (2023), and BLAGN detected with NIRSpec prism spectroscopy at from Greene et al. (2024). Notably, the for those BLAGN were derived using the same methodology employed in this study, whereas Yang et al. (2021) estimated based on the continuum luminosity at 3000 Å and the FWHM of the Mg II line.
@@
The majority of our sample is found to be accreting at an Eddington ratio of the order of 0.1, lower than that of UV-selected luminous quasars, consistent with the intrinsically faint nature of our broad H emitters. These moderate accretion rates are in line with those reported for BLAGN in Maiolino et al. (2024) and Greene et al. (2024). In contrast, the bolometric luminosities of the BLAGN in Harikane et al. (2023) are derived using a more complex methodology, where as estimated from Equation 6 is treated as a lower limit. Taking this into account, we find no significant difference in the inferred accretion rates between our sample and that of Harikane et al. (2023).
@@
Interestingly, we notice that the BLAGN identified with NIRSpec prism() spectroscopy in Greene et al. (2024) tend to occupy the high-mass end of the distribution relative to our sample. The recent JWST Cycle 3 General Observer program Slitless Areal Pure-Parallel High-Redshift Emission Survey (SAPPHIRES-PID: 6434; PI: E. Egami; Sun et al. 2025a) observed one BLAGN (FWHM 4000 km/s) from Greene et al. (2024) with the F444W grism, providing an independent H line measurement. While the source still exhibits a broad feature, the line width appears significantly narrower (FWHM 1300 km/s). Given the grism’s higher spectral resolution, we adopt the latter measurement and mark the resulting and with an “X” in Figure 6, which suggests the possibility that some of the line widths measured by NIRSpec/prism may be overestimated. The details about the analysis of the SAPPHIRES spectrum are presented in Appendix B
V.3 The - Relation
@@
In this work, the stellar masses () of the AGN hosts are measured by SED-fitting using the Prospector Bayesian inference framework (Leja et al. 2017; Johnson et al. 2021, 2022) based on the Flexible Stellar Population Synthesis (FSPS; Conroy et al. 2009; Conroy and Gunn 2010) code. To achieve reliable AGN identification, we adopt a modified version of Prospector presented in Lyu et al. (2024), which incorporates a semi-empirical model for AGN UV-to-mid-IR continua with nebular emission lines and dust attenuation and replaces the default AGN torus model in Prospector.
During the SED-fitting process, we include not only the deep JWST/NIRCam photometry but also multi-band data spanning the optical to near-infrared from HST/ACS and HST/WFC3, as well as longer-wavelength infrared data from Spitzer, if they are available. Examples of galaxy SED-fitting results are shown in Figure 7. The bottom panel presents the only source with a robust detection at longer wavelengths from Spitzer. For the remaining sources, due to the limited wavelength coverage of the NIRCam photometry, clear AGN signatures are identified through SED analysis in only 9 out of 18 galaxies, despite the presence of broad H emission lines. Therefore, follow-up MIRI observations to obtain mid-infrared data for these broad H emitters would be valuable for better constraining their physical properties. Table 4 lists the best-fit values and corresponding uncertainties of for all broad H emitters. Table 4 also summarizes the key physical properties of the 19 broad H emitters.
@@
Figure 8 illustrates the distribution of broad H emitters identified in this study on the plane. To further explore the evolution of the relation over cosmic time, we incorporate AGN recently discovered with JWST across a wide redshift range. These include one potential AGN at , UHZ1 (Bogdán et al., 2024), as well as BLAGN in the GOODS-N field at (Maiolino et al., 2024) and (Sun et al., 2025b), and BLAGN at identified by Harikane et al. (2023). The local relations for AGN hosts derived in Reines and Volonteri (2015) (hereafter RV15) and Greene et al. (2020) (all, limits case, hereafter GSH20) are also included for comparison purposes.
@@
We find that nearly all AGN recently revealed by JWST observations at lie above RV15, with the majority located significantly above the relation. When compared to the GSH20, which is considered a more robust local relation that shows no bias with galaxy type and provides a better match to broad-line selected AGN at (Sun et al., 2025c), the offset is somewhat reduced. Nevertheless, many sources still exhibit higher / ratios than those predicted by GSH20. The over-massive relative to the of host galaxy, when compared with the local scaling relations, is consistent with findings from numerous previous studies of AGN at similar redshifts (e.g., Kocevski et al. 2023; Übler et al. 2023). These results may suggest a different evolution path of the – relation in the early Universe compared to that observed locally. Additionally, the presence of SMBHs with above the local scaling relations aligns with theoretical predictions from models invoking heavy black hole seeds and super-Eddington accretion, which are proposed to explain the rapid growth of SMBHs at very early time (Trinca et al., 2022; Volonteri et al., 2023; Schneider et al., 2023). Alternatively, the growth of the host galaxy can also be suppressed by strong feedback processes (see Zhu et al. 2025 and references therein). However, it is also important to note that, although our AGN sample is selected based on the presence of broad emission lines in spectroscopy, selection bias still exists to some extent. Specifically, due to the detection limits of the grism observations and the requirement of FWHM 1000 km/s, we may fail to detect very faint, low-mass AGN (see Section V.6 for more details), which could otherwise help fill the observed gap between our sample and the local - relation.
V.4 Are Broad Line Emitters LRDs?